# The finite groups with three automorphism orbits

### v1

@@ 5. Subfield hyperplanes and quotients of $q^{1+2m}_{+}$
If (1) holds, then $N$ is an extraspecial $p$-group of exponent $p$, which is in the 7-th row. If (2) holds, then $N$ is in the 6-th row by Lemmas 4.1 and 4.3. Assume that (3) holds. Then $N$ is a quotient of $q^{1+2m}_{+}/U$ for some $U\leqslant{\mathbf{Z}}(q^{1+2m}_{+})$ by Lemma 4.4 (ii), and $q^{1+2m}_{+}$ is listed in the 7-th row. If $U\neq 1$ and contains some subfield hyperplane $U_{0}$, then $N$ is a quotient of $q^{1+2m}_{+}/U_{0}\cong(q_{0})^{(1+2mn_{0})}_{+}$ by Theorem C (i). Thus, we may assume that $U$ contains no subfield hyperplanes. Then, Lemma 5.4 deduces the last row of Table 1, and we complete the proof. ∎
@@
Let $N_{1}:=N/U$, and let $V:=N/{\mathbf{Z}}(N)=N_{1}/{\mathbf{Z}}(N_{1})$. Then ${\mathrm{Aut}}(N)^{V}\cong{\mathrm{C}\Gamma\mathrm{Sp}}(2m,q)$ by Lemma 4.4 (iii). Since ${\mathrm{Sp}}(2m,q)\leqslant{\mathrm{Aut}}(N)^{V}$ acts trivially on ${\mathbf{Z}}(N)$, it follows that ${\mathrm{Sp}}(2m,q)\leqslant{\mathrm{Aut}}(N_{1})^{V}$ and ${\mathrm{Aut}}(N_{1})^{V}$ acts transitively on non-zero vectors of $V$. Thus, $N_{1}$ is $3$-orbit if and only if ${\mathrm{Aut}}(N_{1})^{{\mathbf{Z}}(N_{1})}$ is transitive on non-identity elements of ${\mathbf{Z}}(N_{1})={\mathbf{Z}}(N)/U$ deduced by Lemma 2.8 (iii). It suffices to prove that $(H_{U})^{{\mathbf{Z}}(N_{1})}={\mathrm{Aut}}(N_{1})^{{\mathbf{Z}}(N_{1})}$.
Let $U_{0}$ be a subfield hyperplane with respect to subfield ${\mathbb{F}}_{p^{d}}$ for some $d\mid n$ such that $U_{0}\leqslant U$ and $U_{1}:=U/U_{0}$ has no subfield hyperplanes of ${\mathbf{Z}}(N)/U_{0}\cong{\mathbb{F}}_{p^{d}}^{+}$. By Theorem C (i) and Lemma 4.4 (iii), we have
$N_{0}:=N/U_{0}\cong(q_{0})^{1+2mn/d}_{+}\mbox{ and }H_{0}:={\mathrm{Aut}}(N_{0})^{{\mathbf{Z}}(N)/U_{0}}\cong{\Gamma\mathrm{L}}(1,q_{0}).$
@@
Since $U_{1}$ has no subfield hyperplanes, Lemma 5.4 deduces that
@@
$U_{1}:=\langle 1+\lambda^{2},2\lambda+\lambda^{2}+2\lambda^{3}\rangle\mbox{ and }U_{2}:=\langle 2+\lambda,1+\lambda+\lambda^{2}\rangle.$
We remark that each of $U_{1}$ and $U_{2}$ contains no subfield hyperplanes, and $H_{U_{i}}\cong{\mathbb{Z}}_{8}$ acts transitively on non-zero vectors of $V/U_{i}$ for $i=1,2$ where $H={\Gamma\mathrm{L}}(1,3^{4})$. Hence, $U_{1}$ and $U_{2}$ admit two smallest $3$-orbit groups (of order $3^{10}$) in the 8-th row of Table 1, which are $3^{1+2}_{+}/U_{1}$ and $3^{1+2}_{+}/U_{2}$, respectively. Notice that there exists $h\in N_{H}(L)$ such that $U_{1}^{h}=U_{2}$, the above two groups are isomorphic. We present a generating relation for the above group blow:
@@
Let $p$ and $r$ be distinct odd primes such that $r\nmid p-1$, and let $q=p^{n}$ with $n=p^{r}-1$. Set $V:={\mathbb{F}}_{q}^{+}$, and let $H={\Gamma\mathrm{L}}(1,q)$ acting naturally on $V$. Let $y\in H$ admitted by the Frobenius automorphism in ${\mathrm{Aut}}({\mathbb{F}}_{q})$. Then there exists an $r$-dimensional faithful ${\mathbb{F}}_{p}\langle y\rangle$-submodule $R$ of $V$. Let $U$ be an ${\mathbb{F}}_{p}L$-complement of $R$ in $V$.
With definitions in the above example, we observe that $|V/U|-1=|R|-1=p^{r}-1=|y|$. Hence, $\langle y\rangle$ acts transitively on non-zero vectors of $V/U$, and then $q^{1+2m}_{+}/U$ is a $3$-orbit group for any positive integer $m$ by Corollary 1.1. Since $\dim R=r$ is not a divisor of $\dim V=p^{r}-1$, $q^{1+2m}_{+}/U$ lies in 8-th row of Table 1. Remark that the minimum integer $n$ satisfying the conditions in Example 5.5 is $n:=p^{r}-1=5^{3}-1=124$. Thereofore, the smallest $3$-orbit group admitted by Example 5.5 has order $5^{251}$ with center of order $5^{3}$.

### v2

@@ 5. Subfield hyperplanes and quotients of $q^{1+2m}_{+}$
If (a) holds, then $N$ is an extraspecial $p$-group of exponent $p$, which is in the 7-th row. If (b) holds, then $N$ is in the 6-th row by Lemmas 4.1 and 4.3. Assume that (c) holds. Then $N$ is a quotient of $q^{1+2m}_{+}/U$ for some $U\leqslant{\mathbf{Z}}(q^{1+2m}_{+})$ by Lemma 4.4 (ii), and $q^{1+2m}_{+}$ is listed in the 7-th row. If $U\neq 1$ and contains some subfield hyperplane $U_{0}$, then $N$ is a quotient of $q^{1+2m}_{+}/U_{0}\cong(q_{0})^{(1+2mn_{0})}_{+}$ by Theorem C (i). Thus, we may assume that $U$ contains no subfield hyperplanes. Therefore, Lemma 5.4 implies the last row of Table 1, and we complete the proof. ∎
@@
Let $N_{1}=N/U$, and let $V=N/{\mathbf{Z}}(N)=N_{1}/{\mathbf{Z}}(N_{1})$. Then ${\mathrm{Aut}}(N)^{V}\cong{\mathrm{C}\Gamma\mathrm{Sp}}(2m,q)$ by Lemma 4.4 (iii). Since ${\mathrm{Sp}}(2m,q)\leqslant{\mathrm{Aut}}(N)^{V}$ acts trivially on ${\mathbf{Z}}(N)$, it follows that ${\mathrm{Sp}}(2m,q)\leqslant{\mathrm{Aut}}(N_{1})^{V}$ and ${\mathrm{Aut}}(N_{1})^{V}$ acts transitively on non-zero vectors of $V$. Thus, $N_{1}$ is $3$-orbit if and only if ${\mathrm{Aut}}(N_{1})^{{\mathbf{Z}}(N_{1})}$ is transitive on non-identity elements of ${\mathbf{Z}}(N_{1})={\mathbf{Z}}(N)/U$ by Lemma 2.8 (iii). It suffices to prove that $(H_{U})^{{\mathbf{Z}}(N_{1})}={\mathrm{Aut}}(N_{1})^{{\mathbf{Z}}(N_{1})}$.
Let $U_{0}$ be a subfield hyperplane with respect to subfield ${\mathbb{F}}_{p^{d}}$ for some $d\mid n$ such that $U_{0}\leqslant U$ and $U_{1}=U/U_{0}$ has no subfield hyperplanes of ${\mathbf{Z}}(N)/U_{0}\cong{\mathbb{F}}_{p^{d}}^{+}$. Let $N_{0}=N/U_{0}$, and let $H_{0}={\mathrm{Aut}}(N_{0})^{{\mathbf{Z}}(N)/U_{0}}$. By Theorem C (i) and Lemma 4.4 (iii), we have
$N_{0}\cong(q_{0})^{1+2mn/d}_{+}\mbox{ and }H_{0}\cong{\Gamma\mathrm{L}}(1,q_{0}).$
@@
Since $U_{1}$ has no subfield hyperplanes, Lemma 5.4 implies that
@@
$U_{1}=\langle 1+\lambda^{2},2\lambda+\lambda^{2}+2\lambda^{3}\rangle\mbox{ and}U_{2}=\langle 2+\lambda,1+\lambda+\lambda^{2}\rangle.$
We remark that each of $U_{1}$ and $U_{2}$ contains no subfield hyperplanes, and $H_{U_{i}}\cong{\mathbb{Z}}_{8}$ acts transitively on non-zero vectors of $V/U_{i}$ for $i=1,2$ where $H={\Gamma\mathrm{L}}(1,3^{4})$. Hence, $U_{1}$ and $U_{2}$ admit two smallest $3$-orbit groups (of order $3^{10}$) in the 8-th row of Table 1, which are $3^{1+2}_{+}/U_{1}$ and $3^{1+2}_{+}/U_{2}$, respectively. Notice that there exists $h\in N_{H}(L)$ such that $U_{1}^{h}=U_{2}$, the above two groups are isomorphic. We present a generating relation for the above group below:
@@
Let $p$ and $r$ be distinct odd primes such that $r\nmid p-1$, and let $q=p^{n}$ with $n=p^{r}-1$. Set $V={\mathbb{F}}_{q}^{+}$, and let $H={\Gamma\mathrm{L}}(1,q)$ acting naturally on $V$. Let $y\in H$ admitted by the Frobenius automorphism in ${\mathrm{Aut}}({\mathbb{F}}_{q})$. Then there exists an $r$-dimensional faithful ${\mathbb{F}}_{p}\langle y\rangle$-submodule $R$ of $V$. Let $U$ be an ${\mathbb{F}}_{p}L$-complement of $R$ in $V$.
With definitions in the above example, we observe that $|V/U|-1=|R|-1=p^{r}-1=|y|$. Hence, $\langle y\rangle$ acts transitively on non-zero vectors of $V/U$, and then $q^{1+2m}_{+}/U$ is a $3$-orbit group for any positive integer $m$ by Corollary 1.1. Since $\dim R=r$ is not a divisor of $\dim V=p^{r}-1$, $q^{1+2m}_{+}/U$ lies in 8-th row of Table 1. Remark that the minimum integer $n$ satisfying the conditions in Example 5.5 is $n=p^{r}-1=5^{3}-1=124$. Therefore, the smallest $3$-orbit group admitted by Example 5.5 has order $5^{251}$ with center of order $5^{3}$.